# On the singular loci of higher secant varieties of Veronese embeddings

### v2

@@ Abstract.
In this paper we study singular loci of higher secant varieties of the image of the $d$-uple Veronese embedding of projective $n$-space, $v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N={n+d\choose d}-1}$. We call the closure of the union of all the $(k-1)$-planes spanned by $k$ points on $X$ the $k$-th secant of $X$ for a given embedded variety $X\subset{\mathbb{P}}^{N}$ and denote it by $\sigma_{k}(X)$. For the singular loci of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, it has been known only for $k\leq 3$. As investigating geometry of moving tangents along subvarieties, we determine the (non-)singularity of so-called the $m$-subsecant locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, consisting of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ with any $m$-plane ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$. These loci with arbitrary $k,d,m,n$ show an interesting trichotomy between generic smoothness, non-trivial singularity and trivial singularity. In many cases, they can be served as a new source for the singular locus of the $k$-th secant other than a trivial one, the $(k-1)$-th secant variety. We also consider the case of the $4$-th secant of $v_{d}({\mathbb{P}}^{n})$ by applying main results and computing conormal space via a certain type of Young flattening. In the end, some generalization and discussions for further development are presented as well.
@@
singular locus of higher secant varieties, the subsecant loci, Veronese embedding, symmetric tensors, identifiability
@@ 1. Introduction
$\sigma_{k}(X)=\overline{\bigcup_{x_{1},\dots,x_{k}\in X}\langle\,x_{1},\dots,x_{k}\,\rangle}\subset{\mathbb{P}}^{N}~{},$
where $\langle\,x_{1},\dots,x_{k}\,\rangle\subset{\mathbb{P}}^{N}$ denotes the linear span of the points $x_{1},\dots,x_{k}$ and the overline means the Zariski closure. In particular, $\sigma_{1}(X)=X$ and $\sigma_{2}(X)$ has been often called just the secant or secant line variety of $X$ in the literature.
The secant variety construction (more generally, join construction of subvarieties) is not only one of the most famous methods in classical algebraic geometry, but also a very popular object to be considered in recent years, especially in connection with fields of research such as tensor rank and decomposition, algebraic statistics, data science, geometric complexity theory and so on (see [21, 22] for more details).
Despite of a rather long history and its popularity, most of fundamental questions on the higher secant varieties $\sigma_{k}(X)$ still remain open even for many well-known base variety $X$. For instance, the dimension theorem for $\sigma_{k}(X)$ has been only settled completely for $X$ being curves, Veronese varieties or few other cases. Questions about defining equations of $k$-secant variety also have been answered only for very small $k$’s of few cases and seem far from understanding the essence of sources for the equations (see also [21, chapter 5] for a reference).
In this paper we concentrate on the case of $X=v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$, the Veronese variety, which is the image of the Veronese embedding $v_{d}:{\mathbb{P}}^{n}\hookrightarrow{\mathbb{P}}^{N}$ with $N={n+d\choose d}-1$. In particular, we study the singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, arbitrary higher secant variety of the Veronese variety.
Knowledge on singularities of higher secant varieties is fundamental and very important for its own sake in the study of algebraic geometry and also can be useful for problems in applications. For example, this can be served as a key condition to determine identifiability of structured tensors (see e.g. the introduction in [8] and references therein).
@@
unless $\sigma_{k}(X)$ fills the whole linear span $\langle\,X\,\rangle$. In the paper, we call a point $p\in\sigma_{k}(X)$ a non-trivial singular point if $p\notin\sigma_{k-1}(X)$ and $\sigma_{k}(X)$ is singular at $p$, while the points belonging to $\sigma_{k-1}(X)$ are called trivial singular points.
@@
There have been some known results on the singular loci of $\sigma_{k}(X)$ mostly for very small $k$’s. The equality in (1) holds for determinantal varieties defined by minors of a generic matrix. It is completely described for the secant variety of the Segre product of projective spaces (i.e., $k=2$) in [27]. It has been recently generalized to the case of $\sigma_{2}(X)$ where $X$ is a Segre-Veronese embedding by [2] and $X$ is a Grassmannian by [13]. For the $3$-th secant variety of Grassmannian $\mathbb{G}(2,6)$, dimension and some other properties of the singular loci was studied in [4].
@@
For the case of Veronese varieties, it is classical that $\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))=\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ holds both for the binary case (i.e., $n=1$) and for the case of quadratic forms (i.e., $d=2$) (see e.g. [19, chapter 1]). When $k=2$, it was proved that the equality above (i.e., having only trivial singular locus) is also true for any $d,n$ by [20]. For $k=3$, by the second author the singular locus was completely determined in [16]. In particular, it was shown that a non-trivial singularity can happen and it is only when $d=4$ and $n\geq 3$.
@@
In the present paper, we explore on the singular loci of any higher secant varieties of Veronese varieties and introduce another main origin for the singularity of $k$-secant variety other than the trivial singular locus. We call this the ‘subsecant locus’ and it shows an interesting trichotomy phenomenon among generic smoothness, non-trivial singular points and trivial singularity as in our main results.
@@
For any given point $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, as considering $(1,d-1)$-symmetric flattening (see §3 for details), it is immediate to see that there exists an $m$-plane (i.e., $m$-dimensional linear subvariety) ${\mathbb{P}}^{m}$ of ${\mathbb{P}}^{n}$ with $11\leq m\leq k-$ 1 such that $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$. So, from now on, we call $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ an $m$-subsecant variety of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ (or simply a subsecant variety in case of no confusion) if $m<k-1$ and $m<n$. We also define
(2) $\Sigma_{k,d}(m)=\Sigma_{k,d}(m;{{\mathbb{P}}^{n}}):=\bigcup_{\forall{\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}}\sigma_{k}(v_{d}({\mathbb{P}}^{m})),$
the $m$-subsecant locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. It naturally forms an increasing sequence of loci in the $k$-secant as
$\Sigma_{k,d}(1)\subset\Sigma_{k,d}(2)\subset\cdots\subset\Sigma_{k,d}(\min\{k-1,n\}-1)\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))=\Sigma_{k,d}(\min\{k-1,n\})~{}.$
@@
We call $\Sigma_{k,d}(\min\{k-1,n\}-1)$ (i.e., the union of all proper subsecant varieties) the maximum subsecant locus of the given $k$-th secant $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Any point outside the maximum subsecant locus and the previous secant variety (i.e., $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))\setminus\big{(}\Sigma_{k,d}(\min\{k-1,n\}-1)\cup\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))\big{)}$) is called a point of full-secant locus. Note that in the case of $k=3$ of [16], the only exceptional case for $\operatorname{Sing}(\sigma_{3}(v_{d}({\mathbb{P}}^{n})))$ (i.e., $d=4$ and $n\geq 3$), the singular locus is given as the maximum subsecant locus $\Sigma_{3,d}(1)$, while the locus is just the trivial one, $\sigma_{2}(v_{d}({\mathbb{P}}^{n}))$, for all the other $d,n$ (see Remark 4 (a)). Most of previously known results on singular loci of secant varieties can be understood in this viewpoint (see Remark 37).
@@
When does an $m$-subsecant locus $\Sigma_{k,d}(m)$ become singular in $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$?
In principle, it is in some sense (in spite of its computational complexity) straightforward to check the singularity, once a complete set of equations for a higher secant variety is attained. But, as mentioned above, not much has been known on the defining equations and it seems far from being understood completely at this moment, even for the Veronese case (see [23, 15] for the state of the art). Lacking in knowledge on the equations for higher secant varieties, it is very difficult to determine singular locus in general.
In this paper, without further understanding on equations (!), we introduce a geometric way to pursue it for this kind of problems, which is based on a careful study on behavior of embedded tangent spaces moving along a locus in the Veronese variety . For $m=1$ case, we first present the following result for the $1$-subsecant locus $\Sigma_{k,d}(1)=\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, which is a generalization of [16, theorem 2.1] (i.e., $k=3$ case) to any higher $k$-secant of Veronese varieties.
@@
    If $k=\frac{d+2}{2}$ , then $\Sigma_{k,d}(1)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(1)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity). This case happens only if $d$ is even.
@@
    If $k\geq\frac{d+3}{2}$, then $\Sigma_{k,d}(1)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity).
@@
Concerning singular points of arbitrary $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ originated from subsecant loci, we prove the following general theorems on the $m$-subsecant locus $\Sigma_{k,d}(m)$ with $m\geq 2$ and $k\geq 4$ as our main results.
@@
Let $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ be the $d$-uple Veronese embedding with $n\geq 3$, $d\geq 3$, and $N={n+d\choose d}-1$. Let $k\geq 4$ and let ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ be an $m$-plane with $2\leq m<\min\{k-1,n\}$. Assume that $(d,m)\notin\{(3,4),(4,2),(4,3),(4,4)\}$.
For $(k,d,n)\neq(4,3,3)$, setting $\mu$ to be ($\frac{{m+d\choose m}}{m+1}$ if $(d,m)\notin\{(6,2),(3,5)\}$ and to be ($\frac{{m+d\choose m}}{m+1}-1$ if $(d,m)\in\{(6,2),(3,5)\}$, we have:
@@
    If $k\leq\mu$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at a general point in $\Sigma_{k,d}(m)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
@@
    If $\mu<k<\mu+1$, then $\Sigma_{k,d}(m)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(m)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity). In particular, this singularity appears if and only if $\mu\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$.
@@
    If $k\geq\mu+1$, then $\Sigma_{k,d}(m)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity).
@@
For the cases of $(d,m)=(3,4),(4,2),(4,3)$ or $(4,4)$, only in which the dimension of a $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ can be defective, we have the following statement.
@@
In the same situation as Theorem 2, for $(d,m)\in\{(3,4),(4,2),(4,3),(4,4)\}$ and $\mu=\frac{{m+d\choose m}}{m+1}$, the following holds:
@@
    If $k\leq\lfloor\mu\rfloor-1$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at a general point in $\Sigma_{k,d}(m)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
@@
    If $\lfloor\mu\rfloor\leq k\leq\lceil\mu\rceil+1$, then $\Sigma_{k,d}(m)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(m)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
@@
    If $k\geq\lceil\mu\rceil+2$, then $\Sigma_{k,d}(m)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
We summarize Theorem 3 in Table 1 with exact ranges of $k$.
$(d,m)$ $\Sigma_{k,d}(m)$ : Generically smooth (i) Non-trivial singular locus (ii) Trivial singularity (iii)
$(3,4)$ $k\leq 6$ $k=7,8$ $9\leq k$
$(4,2)$ $k\leq 4$ $k=5,6$ $7\leq k$
$(4,3)$ $k\leq 7$ $k=8,9,10$ $11\leq k$
$(4,4)$ $k\leq 13$ $k=14,15$ $16\leq k$
Table 1. (Non-)singularity of $\Sigma_{k,d}(m)$ in case of defective $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s in Theorem 3
@@
We’d like to make some remarks on the theorems above.
@@
    In the case of $k=3$ and $n\geq 3$, the condition of (ii) in Theorem 1 holds if and only if $d=4$, and then $\operatorname{Sing}(\sigma_{3}(v_{4}({\mathbb{P}}^{n}))$ is equal to $\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\langle\,v_{4}({\mathbb{P}}^{1})\,\rangle=\Sigma_{3,4}(1)$ since $\sigma_{3}(v_{4}({\mathbb{P}}^{1}))=\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle$ and $\sigma_{2}(v_{4}({\mathbb{P}}^{n}))\subset\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\langle\,v_{d}({\mathbb{P}}^{1})\,\rangle$ (see also Corollary 32). This gives a geometric description of the only exceptional case for the singular loci of $3$rd secant variety in [16].

### v3

@@ Abstract.
The $k$-th secant variety of a projective variety $X\subset\mathbb{P}^{N}$, denoted by $\sigma_{k}(X)$, is defined to be the closure of the union of $(k-1)$-planes spanned by $k$ points on $X$. In this paper, we examine the $k$-th secant variety $\sigma_{k}(v_{d}(\mathbb{P}^{n}))\subset\mathbb{P}^{N}$ of the image of the $d$-uple Veronese embedding $v_{d}$ of $\mathbb{P}^{n}$ to $\mathbb{P}^{N}$ with $N=\binom{n+d}{d}-1$, and focus on the singular locus of $\sigma_{k}(v_{d}(\mathbb{P}^{n}))$, which is only known for $k\leq 3$. To study the singularity for arbitrary $k,d,n$, we define the $m$-subsecant locus of $\sigma_{k}(v_{d}(\mathbb{P}^{n}))$ to be the union of $\sigma_{k}(v_{d}(\mathbb{P}^{m}))$ with any $m$-plane $\mathbb{P}^{m}\subset\mathbb{P}^{n}$. By investigating the projective geometry of moving embedded tangent spaces along subvarieties and using known results on the secant defectivity and the identifiability of symmetric tensors, we determine whether the $m$-subsecant locus is contained in the singular locus of $\sigma_{k}(v_{d}(\mathbb{P}^{n}))$ or not. Depending on the value of $k$, these subsecant loci show an interesting trichotomy between generic smoothness, non-trivial singularity, and trivial singularity. In many cases, they can be used as a new source for the singularity of the $k$-th secant variety of $v_{d}(\mathbb{P}^{n})$ other than the trivial one, the $(k-1)$-th secant variety of $v_{d}(\mathbb{P}^{n})$. We also consider the case of the $4$-th secant variety of $v_{d}(\mathbb{P}^{n})$ by applying main results and computing conormal space via a certain type of Young flattening. Finally, we present some generalizations and discussions for further developments.
@@
singular locus of higher secant varieties, the subsecant loci, Veronese embedding, symmetric tensors, secant defectivity, identifiability
@@ 1. Introduction
(1) $\sigma_{k}(X)=\overline{\bigcup_{x_{1},\dots,x_{k}\in X}\langle\,x_{1},\dots,x_{k}\,\rangle}\subset{\mathbb{P}}^{N},$
where $\langle\,x_{1},\dots,x_{k}\,\rangle\subset{\mathbb{P}}^{N}$ denotes the linear span of the points $x_{1},\dots,x_{k}$ and the overline means the Zariski closure. In particular, $\sigma_{1}(X)=X$ and $\sigma_{2}(X)$ is often simply called the secant or secant line variety of $X$ in the literature.
The construction of secant varieties (or more generally, join construction of subvarieties) is not only one of the most famous methods in classical algebraic geometry, but also a very popular subject in recent years, especially in connection with fields of research such as tensor rank and decomposition, algebraic statistics, data science, geometric complexity theory and so on (see [Land12, Land17] for more details).
Despite of a rather long history and the popularity, most of the fundamental questions on the higher secant varieties $\sigma_{k}(X)$ still remain open even for many well-known base varieties $X$. For instance, one can ask the secant defectivity question, which concerns the dimension of $\sigma_{k}(X)$. We say that $\sigma_{k}(X)$ is secant defective (or simply defective) if the dimension of $\sigma_{k}(X)$ is less than the expected one, $\min\{N,k\cdot\dim(X)+k-1\}$. It is classically known that higher secant varieties of curves are non-defective (e.g. [Rus16, Corollary 1.2.3]). Due to the famous theorem of Alexander-Hirschowitz [AH95], we know the dimensions of higher secant varieties of all Veronese varieties. In other research, there are only a few cases where the dimension theorem for $\sigma_{k}(X)$ is fully determined (see [Rus16], [Zak93] for references). Questions about defining equations of $k$-th secant varieties have also only been answered for very small $k$’s of a few cases and seem still far from understanding the essence of the sources for the equations (see also [Land12, Chapter 5] for a reference).
In this paper, we concentrate on the case of $X=v_{d}({\mathbb{P}}^{n})\subset{\mathbb{P}}^{N}$, the Veronese variety, which is the image of the Veronese embedding $v_{d}:{\mathbb{P}}^{n}\hookrightarrow{\mathbb{P}}^{N}$ with $N={n+d\choose d}-1$. In particular, we study the singular locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, an arbitrary higher secant variety of the Veronese variety.
The knowledge on singularities of higher secant varieties is fundamental and very important for its own sake in the study of algebraic geometry and also can be useful for problems in applications. For example, it can be used as a key condition to establish the identifiability of structured tensors (e.g. the introduction in [COV14] and references therein).
@@
unless $\sigma_{k}(X)$ fills up the whole linear span $\langle\,X\,\rangle$ (see [Rus16, Proposition 1.2.2]). In the paper, we say that a point $p\in\sigma_{k}(X)$ is a non-trivial singular point if $p\notin\sigma_{k-1}(X)$ and $\sigma_{k}(X)$ is singular at $p$, while the points belonging to $\sigma_{k-1}(X)$ are called trivial singular points.
@@
There are some known results on the singular loci of $k$-th secant varieties $\sigma_{k}(X)$, mostly for very small $k$’s. The equality in (2) holds for determinantal varieties defined by minors of a generic matrix. It is completely described for the $2$-nd secant variety of the Segre product of projective spaces in [MOZ15]. It has recently been generalized to the case of $\sigma_{2}(X)$, where $X$ is a Segre-Veronese embedding by [AMZ20] and $X$ is a Grassmannian by [GS24]. For the $3$-rd secant variety of the Grassmannian $\mathbb{G}(2,6)$, the dimension and some other properties of the singularity have been studied in [AOP12].
@@
For the case of Veronese varieties, it is classical that $\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))=\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ holds both for the binary case (i.e., $n=1$) and for the case of quadratic forms (i.e., $d=2$) (see e.g. [IK99, Chapter 1]). For $k=2$, it was proved in [Kan99] that the above equality holds also for any $d,n$. In these cases, the $k$-th secant variety has only the trivial singularity. For $k=3$, the singular locus was completely determined by the second author in [Han18]; in particular, the non-trivial singularity occur if and only if $d=4$ and $n\geq 3$.
@@
In the present paper, we explore the singular locus of any higher secant variety of the Veronese variety, and introduce a new main origin for the singularity other than the trivial singularity. We call this the ‘subsecant locus’. As in our main results, these loci show an interesting trichotomy phenomenon among generic smoothness, non-trivial singularity, and trivial singularity.
@@
For any given point $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, there exists an $m$-plane (i.e., $m$-dimensional linear subvariety) ${\mathbb{P}}^{m}$ of ${\mathbb{P}}^{n}$ with $11\leq m\leq k-$ 1 such that $p\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$; it immediately follows for a general $p$, and even if $p$ is in the boundary of the closure in (1), it is also true by considering $(1,d-1)$-symmetric flattening (see §3 for details). So, from now on, we say that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is an $m$-subsecant variety of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ if $m<k-1$ and $m<n$, and simply call it a subsecant variety in case there is no confusion. We also define the $m$-subsecant locus of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$,
[…]
It naturally forms an increasing sequence of loci in the $k$-th secant variety as
$\Sigma_{k,d}(1)\subset\Sigma_{k,d}(2)\subset\cdots\subset\Sigma_{k,d}(\min\{k-1,n\}-1)\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))=\Sigma_{k,d}(\min\{k-1,n\}).$
@@
In particular, $\Sigma_{k,d}(\min\{k-1,n\}-1)$ is the union of all proper subsecant varieties, which we call the maximum subsecant locus of the given $k$-th secant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Any point of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ outside the maximum subsecant locus and the previous secant variety $\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ is called a point of the full-secant locus. Note that for $k=3$ of [Han18], when $d=4$ and $n\geq 3$, the singular locus of $\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ is given as the maximum subsecant locus $\Sigma_{3,d}(1)$, which is the only case where the singularity pattern of $3$-rd secant varieties becomes exceptional, while for all the other $d,n$ the singularity is just the trivial one, $\sigma_{2}(v_{d}({\mathbb{P}}^{n}))$ (see Remark 4 (a)). Most of the previously known results on singular loci of secant varieties can be understood in this viewpoint (see Remark 37).
@@
When is $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ singular at points of an $m$-subsecant locus $\Sigma_{k,d}(m)$?
In principle, it is somewhat straightforward (despite the computational complexity) to check the singularity, once a complete set of equations for a higher secant variety is attained. But, as mentioned above, not much is known about the defining equations and they seem quite far from being fully understood at this moment, even for the Veronese case (see [LO13, FH24] for the state of the art). Due to the lack of knowledge on the equations for the higher secant variety, it is very difficult to determine the singular locus in general.
In this paper, without further understanding on the equations (!), we introduce a geometric way to pursue it for this kind of problems, which is based on a careful study on the behavior of embedded tangent spaces moving along a locus in the Veronese variety . For the case $m=1$, we first present the following result for the $1$-subsecant locus $\Sigma_{k,d}(1)=\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ of $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, which is a generalization of [Han18, Theorem 2.1] (i.e., $k=3$ case) to any higher $k$-th secant varieties of Veronese varieties.
@@
    If $k=\frac{d+2}{2}$ , then $\Sigma_{k,d}(1)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(1)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity). This case occurs only if $d$ is even.
@@
    If $k\geq\frac{d+3}{2}$, then $\Sigma_{k,d}(1)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity, unless $\sigma_{k}(v_{d}(\mathbb{P}^{n}))=\mathbb{P}^{N}$).
@@
Concerning singular points of arbitrary $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ originated from subsecant loci, we prove the following general theorems on the $m$-subsecant locus $\Sigma_{k,d}(m)$ with $m\geq 2$ and $k\geq 4$ as the main results.
@@
Let $v_{d}:{\mathbb{P}}^{n}\rightarrow{\mathbb{P}}^{N}$ be the $d$-uple Veronese embedding with $n\geq 3$, $d\geq 3$, and $N={n+d\choose d}-1$. Let $k\geq 4$ and let ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ be an $m$-plane with $2\leq m<\min\{k-1,n\}$. Assume that $(d,m)\notin\mathcal{E}=\{(3,3),(3,4),(3,5),(4,2),(4,3),(4,4),(5,2),(6,2)\}$.
For $(k,d,n)\neq(4,3,3)$, setting $\mu=\Bigr{\lceil}\frac{\binom{m+d}{m}}{m+1}\Bigl{\rceil}$, we have:
@@
    If $k<\mu$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at a general point in $\Sigma_{k,d}(m)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
@@
    If $k=\mu$, then $\Sigma_{k,d}(m)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(m)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity).
@@
    If $k>\mu$, then $\Sigma_{k,d}(m)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity, unless $\sigma_{k}(v_{d}(\mathbb{P}^{n}))=\mathbb{P}^{N}$).
@@
In the same situation as Theorem 2, for
$(d,m)\in\mathcal{E}=\{(3,3),(3,4),(3,5),(4,2),(4,3),(4,4),(5,2),(6,2)\},$
if $k$ is in one of the ranges named (i), (ii), (iii) in Table 1 below, then the following property corresponding to the name of the range holds.
@@
    $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is smooth at a general point in $\Sigma_{k,d}(m)\setminus\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
@@
    $\Sigma_{k,d}(m)\subset\operatorname{Sing}(\sigma_{k}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{k,d}(m)\not\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
@@
    $\Sigma_{k,d}(m)\subset\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$.
$(d,m)$ $\frac{\binom{m+d}{m}}{m+1}$ (i) (ii) (iii)
$(3,3)$ $5$ $k\leq 5$ None $6\leq k$
$(3,4)$ $7$ $k\leq 6$ $k=7,8$ $9\leq k$
$(3,5)$ ${28}/{3}$ $k\leq 8$ $k=9,10$ $11\leq k$
$(4,2)$ $5$ $k\leq 4$ $k=5,6$ $7\leq k$
$(4,3)$ ${35}/{4}$ $k\leq 7$ $k=8,9,10$ $11\leq k$
$(4,4)$ $14$ $k\leq 13$ $k=14,15$ $16\leq k$
$(5,2)$ $7$ $k\leq 7$ None $8\leq k$
$(6,2)$ ${28}/{3}$ $k\leq 8$ $k=9,10$ $11\leq k$
Table 1. (Non-)singularity of $\Sigma_{k,d}(m)$ in Theorem 3
To understand the reason for considering the conditions that $(d,m)$ is in $\mathcal{E}$ or not, we discuss the secant defectivity and the generic identifiability of an $m$-subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. Set ${\mathbb{P}}^{\beta}=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle\subset{\mathbb{P}}^{N}$ with $\beta=\binom{m+d}{m}-1$. where $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subset{\mathbb{P}}^{\beta}$ is of dimension $\leq km+k-1$.
From [AH95], the codimension of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ in ${\mathbb{P}}^{\beta}$ is greater than $\max\{\beta-(km+k-1),0\}$ (i.e., $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ does not fill ${\mathbb{P}}^{\beta}$ and is secant defective) if and only if $d=2$ and $2\leq k\leq m$, or $(k,d,m)=(7,3,4),(5,4,2),(9,4,3),(14,4,4)$; indeed, in the latter case, the four $k$-th secant varieties $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s are hypersurfaces of ${\mathbb{P}}^{\beta}$. Except these defective cases, we have $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$ if $km+k-1\geq\beta$, or equivalently $k\geq{\binom{m+d}{m}}/({m+1})$.
We say that a point $a\in{\mathbb{P}}^{\beta}$ is $k$-identifiable if there is a unique $k$-tuple of points $x_{1},\dots,x_{k}\in v_{d}({\mathbb{P}}^{m})$ such that $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$. We also say that $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is generically identifiable if a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is $k$-identifiable. From [GM17, Theorem 1], a general point $a\in{\mathbb{P}}^{\beta}$ is $k$-identifiable (or $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is generically identifiable in the case of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$) if and only if $m=1$ and $k=(d+1)/2$, or $(k,d,m)=(5,3,3),(7,5,2)$. From [COV17, Theorem 1.1], in the case when $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta}$ is not secant defective, $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ is not generically identifiable if and only if $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$.
@@
We make some remarks on the theorems above.
@@
    In the case of $k=3$ and $n\geq 3$, the condition of (ii) in Theorem 1 holds if and only if $d=4$, and then $\operatorname{Sing}(\sigma_{3}(v_{4}({\mathbb{P}}^{n}))$ is equal to $\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\langle\,v_{4}({\mathbb{P}}^{1})\,\rangle=\Sigma_{3,4}(1)$ since $\sigma_{3}(v_{4}({\mathbb{P}}^{1}))=\langle\,v_{4}({\mathbb{P}}^{1})\,\rangle$ and $\sigma_{2}(v_{4}({\mathbb{P}}^{n}))\subset\bigcup_{{\mathbb{P}}^{1}\subset{\mathbb{P}}^{n}}\langle\,v_{4}({\mathbb{P}}^{1})\,\rangle$ (see also Corollary 32). This gives a geometric description of the only exceptional case for the singular loci of the $3$-rd secant varieties in [Han18].